OpenStudy (anonymous):

k(s)=3s^2-s+6 K(-s) ; k(2s)

OpenStudy (bahrom7893):

k(-s) = 3(-s)^2 -(-s) + 6 = 3s^2 + s +6

OpenStudy (bahrom7893):

k(2s) = 3(2s)^2 - (2s) + 6 = 3*4s^2 -2s +6 = 12s^2 - 2s + 6

OpenStudy (anonymous):

so that -(-s) becomesa positive s+6

OpenStudy (bahrom7893):

yup double negative is a positive.

OpenStudy (anonymous):

how does 3(-s)^2 become 3s^2 rather than just 3s ?

OpenStudy (bahrom7893):

anything to an even power becomes positive itself to an even power

OpenStudy (bahrom7893):

(-2)^2 = 2^2 = 4 (-a)^4 = a^4 (t)^2 = t^2, (-t)^2 = t^2

OpenStudy (anonymous):

oh okay..so i dont look at it as (-s)(-s)=s...it just remains squared?

OpenStudy (bahrom7893):

well (-s)(-s) = s^2.. like i said (-s)(-s) = -1(s^1)*-1(s^1) = (-1*-1)(s^1 * s^1) = 1*s^(1+1) = s^2